2) A basketball player scores 70% of his shots on average. What is the probability that he
scores at least 18 successful shots tonight if he gets 20 shots?
Answer #1

n = 20, x = 18, p = 0.7

Using Binomial distribution and Excel's function BINOM.DIST.RANGE(n, p, start x, end x)

Probability that the player scores at least 18 successful shots tonight if he gets 20 shots: P(x > 18)

= P(x=18) + P(x=19) + P(x=20)

= BINOM.DIST.RANGE(20,0.7,18,20)

= 0.03548